OpenStudy (anonymous):

Graph the unit circle using parametric equations with your calculator set to degree mode. Use a scale of 5. Trace the circle to find all values of t between 0° and 360° satisfying the following statement. (Enter your answers as a comma-separated list.) sin t = −cos t

OpenStudy (anonymous):

The equations are \(x(t) = \cos t\) and \(y(t) = \sin t\). As for the solutions to the equation: \[\sin t = -\cos t \\ \frac{\sin t}{\cos t} = -1 \\ \tan t = -1 \\ t = \frac{3\pi}{4}, \frac{7\pi}{4} \]